# Rethinking Diverse Human Preference Learning through Principal Component Analysis

### v1

@@ Abstract
Understanding human preferences is crucial for improving foundation models and building personalized AI systems. However, preferences are inherently diverse and complex, making it difficult for traditional reward models to capture their full range. While fine-grained preference data can help, collecting it is expensive and hard to scale. In this paper, we introduce Decomposed Reward Models (DRMs), a novel approach that extracts diverse human preferences from binary comparisons without requiring fine-grained annotations. Our key insight is to represent human preferences as vectors and analyze them using Principal Component Analysis (PCA). By constructing a dataset of embedding differences between preferred and rejected responses, DRMs identify orthogonal basis vectors that capture distinct aspects of preference. These decomposed rewards can be flexibly combined to align with different user needs, offering an interpretable and scalable alternative to traditional reward models. We demonstrate that DRMs effectively extract meaningful preference dimensions (e.g., helpfulness, safety, humor) and adapt to new users without additional training. Our results highlight DRMs as a powerful framework for personalized and interpretable LLM alignment.
@@
Anonymous ACL submission
@@ 1 Introduction
Reinforcement Learning from Human Feedback (RLHF) stiennon2020learning; ouyang2022training; bai2022training; team2023gemini; grattafiori2024llama3herdmodels; liu2024deepseek has proven to be a powerful approach for fine-tuning large language models (LLMs) to serve as better general assistants team2023gemini; achiam2023gpt, and AI agents nakano2021webgpt; wu2023autogen; zhao2024expel. Typically, LLMs are optimized using a scalar reward model trained on human preference data, acting as a proxy for overall user satisfaction. However, this approach has two key limitations: (1) it often reflects the preferences of the majority, potentially marginalizing underrepresented groups chakraborty2024maxmin; chidambaram2024direct and failing to capture the full diversity of human preferences, and (2) it struggles to represent the complex, multifaceted, and sometimes conflicting nature of human preferences with a single scalar reward jang2023personalized; rame2024rewarded; yang2024rewards; zhou2024beyond.
@@
As demand grows for more personalized LLMs zhang2024personalization, researchers have explored ways to capture fine-grained, multidimensional human preferences. Some studies have introduced datasets that evaluate multiple aspects such as relevance, correctness, completeness, helpfulness, and harmlessness wu2023fine; wang2023helpsteer; cui2023ultrafeedback; pitis2024improving. Building on these datasets, others have proposed multi-objective optimization methods to accommodate diverse user needs yang2024rewards; yang2024metaaligner; wang2024conditional. However, these approaches face significant challenges: collecting fine-grained human annotations is expensive, and using GPT-generated labels can introduce biases. Considering that large-scale binary preference datasets—where users simply compare two responses—are easier to collect and more widely available, we ask the question: Can we infer multidimensional human preferences directly from large-scale binary comparisons?
@@
To address this question, we propose Decomposed Reward Models (DRMs), a framework that extracts fine-grained human preferences using binary comparison data. Unlike traditional probabilistic models such as Bradley-Terry (BT) bradley1952rank or pairwise preference models jiang2023llm, our approach represents human preferences as $d$-dimensional vectors, corresponding to the learned weights of the final linear layer in a reward model. We show that this vector-based representation establishes a connection between Principal Component Analysis (PCA) and preference learning. Our method constructs a dataset of embeddings for all prompt-response pairs and applies PCA to the differences between chosen and rejected response embeddings. This process identifies orthogonal basis vectors, each capturing a distinct human preference direction. These basis vectors can be combined with their corresponding feature extractor to construct reward models. The framework is shown in Figure 1. DRMs are particularly effective for personalized preference learning. Given a small adaptation dataset from a new user, we compute coefficients for the basis vectors and form a linear combination that best aligns with the user’s preferences. This approach allows for flexible, scalable, and interpretable preference modeling without requiring additional training.
@@ 2 Preliminary
The Bradley-Terry (BT) model bradley1952rank is a probabilistic framework commonly used in preference learning. Given a prompt $x$ and two responses, $y_{1}$ and $y_{2}$, the BT model defines the probability of $y_{1}$ being preferred over $y_{2}$ as:
@@ 3.1 Vector Representation of Preferences
In this paper, we consider a vector representation of human preference — denoting $\phi(x,y)\in\mathbb{R}^{d}$ as a $d$-dimensional feature extractor (e.g., the penultimate layer outputs), and a final linear layer $w$ in the reward model, where $r_{\theta}(x,y)=w^{T}\phi(x,y)$. Under this formulation, the BT objective can be rewritten as:
@@ 3.3 From Scalar Reward to Diverse Rewards
Our analysis above suggests a connection between the eigenvectors of the covariance matrix and human preferences. A key observation is that the covariance matrix has $d$ eigenvectors—for example, $d=2048$ for gemma-2B team2024gemma and $d=4096$ for Llama-3.1-8B grattafiori2024llama3herdmodels. This means we can extract a large number of meaningful preference vectors $w$ from PCA applied to the embedding dataset $\{z_{i}\}_{i=1}^{N}$.
@@ 3.4 Decomposed Reward Models (DRMs)
(2) Principal Decomposition: Given the dataset $\mathcal{D}_{e}$ of shape $N\times d$, we perform PCA to obtain a set of basis vectors $W=[w_{1},w_{2},\dots,w_{d}]$. Each eigenvector $w_{i}$ captures a distinct directional human preference and can be combined with the feature extractor $\phi(x,y)$ to construct a reward model, leading to a total of $d$ rewards.
To fully utilize DRMs, we aim to represent any human preference as a linear combination of these decomposed rewards. Various methods can achieve this, such as mixture-of-experts quan2024dmoerm; wang2024interpretable and test-time alignment lee2024test. Since both approaches are fundamentally similar, and test-time alignment can better leverage our method without additional fine-tuning, we adopt it as our implementation. However, DRMs are compatible with any downstream method that requires multiple diverse reward heads.
We implement test-time alignment based on HyRe lee2024test, which adapts to diverse human preferences by optimizing reward weights given a small adaptation set $\mathcal{D}_{\rm adapt}$. This is practical in real-world scenarios where new users’ data helps refine the strategy to better align with individual preferences. The weights $k_{i}$ are computed based on the loss function over the adaptation set:
$k_{i}=\frac{\exp{(-\mathcal{L}(w_{i},\mathcal{D_{\rm adapt}}))}}{\sum_{j}\exp{(-\mathcal{L}(w_{j},\mathcal{D_{\rm adapt}}))}}$ (7)
@@
$\mathcal{L}(w_{i},\mathcal{D_{\rm adapt}})=-\mathbb{E}_{(x_{i},y_{i}^{c},y_{i}^{r})\sim\mathcal{D}_{\rm adapt}}$ (8)
$\left[\log\sigma\left(w^{T}(\phi(x_{i},y_{i}^{c})-\phi(x_{i},y_{i}^{r}))\right)\right]$
Here, we also normalize the features based on $\mathcal{D}_{\rm adapt}$ to prevent the loss $\mathcal{L}$ from being dominated by samples with large-scale features. This formulation ensures that weights $k_{i}$ dynamically adjust based on the dataset, assigning higher importance to directions $w_{i}$ with lower loss and reducing the influence of those with higher loss. As a result, we can obtain a soft, weighted sum of DRMs that better aligns with the desired human preference.
@@
Table 1: Performance of top 100 decomposed reward heads. “Single Head" is the trained single-head baseline. “Max Value" refers to the highest score achieved for each attribute, while “Max Head" indicates which specific head attains this maximum score. “Overall" represents the average accuracy of a single head across all attributes.
@@ 4.1 Experimental Setup
Dataset. We choose the mixture2_and_safe_pku dataset [^1] https://huggingface.co/datasets/weqweasdas/preference_dataset_mixture2_and_safe_pku dong2024rlhf, a collection of 550k pairwise preference samples. This dataset combines diverse data sources, including human-labeled preferences from HH-RLHF bai2022training and GPT-labeled preferences from UltraFeedback cui2023ultrafeedback, making it well-suited for studying diverse preference decomposition. To evaluate the effectiveness of the decomposed reward heads, we test models on two unseen benchmarks with multiple attributes: (1) RewardBench lambert2024rewardbench, a dataset designed to evaluate reward models across various dimensions, including chat quality, safety, and reasoning. (2) Reasonable Preference Reversal (RPR) test set pitis2024improving focuses on personalized context-aware preference evaluation. From RPR, we sample five fine-grained categories (i.e., User-Friendliness, Narrative Quality, Linguistic Creativity, Scientific Rigor, and Humor), each with over 80 annotated samples. Other categories were excluded due to insufficient data for reliable evaluation.
Base Model. For experiments on decomposed reward heads and test-time adaptation, we use two open-source reward models, Gemma-2B-RM and Llama3-8B-RM yang2024regularizing, along with an instruction-tuned language model, gemma-2-9b-it team2024gemma, as our base models. To analyze their performance, we keep the backbone fixed as our feature extractors while generating multiple new reward heads for them.
Baselines. We compare the performance of DRMs against several baselines. (1) Single-Head RM fine-tunes a single reward head on top of a fixed backbone using the same dataset as DRMs. (2) Share-Base RM lee2024test is a training-based method that employs multiple learnable reward heads while incorporating a frozen prior networkosband2023epistemic to maintain diversity, with the final output derived from their combination. (3) Random Head uses multiple reward heads with randomly initialized weights to capture diverse but largely unstructured preferences. Specifically, we experiment with both uniform and Gaussian initialization, followed by L2 normalization to ensure consistency with DRM vectors.
@@
Table 2: Evaluation Results on RPR ($n=5$). We compare DRMs with trained baselines (“Single Head" and “Shared-Base"), and randomly generated multi-head baselines (“Random"). Except for single-head baseline, other methods use HyRe for test-time adaptation. Standard deviation over 20 sampled adaptation sets are reported.
@@
Benchmark Attributes Single Head Random (Uniform) DRM(Ours)
@@ 4.2 What information is Captured by DRMs?
We aim to better understand the decomposed reward heads in DRMs. To achieve this, we evaluate the performance of the top 100 reward vectors, ranked by eigenvalue, on both RewardBench and RPR’s fine-grained subsets. Table 1 reports the scores of a trained single-head baseline. The “Max Value" column shows the highest score achieved for each attribute, while the “Max Head" column indicates which reward head achieves this score. We also compare the results with the single-head baseline. The results reveal the following findings:
(1) Diversity and Interpretability: DRMs effectively capture diverse human preferences, with different reward heads excelling at different attributes. For instance, in Gemma-2B-RM, head_9 performs best on “User-Friendliness" (accuracy: 0.798) and “Humor and Entertainment" (0.964), while head_12 excels in “Narrative and Storytelling" (0.825) and “Linguistic Creativity" (0.885). In contrast, the single-head RM fails to capture this diversity, yielding suboptimal performance on attributes such as “User-Friendliness" (0.506) and “Humor and Entertainment" (0.690) for Gemma-2B-RM. These results indicate that DRMs not only capture a broader range of human preferences but also provide interpretable representations that align well with certain preference attributes.
(2) The first head is the most informative: An interesting observation is that the head head_0 consistently achieves the highest overall accuracy for both models. This aligns with expectations, as head_0 corresponds to the eigenvector with the largest variance, i.e., the most informative direction. Furthermore, among the top 100 heads, most of the high-performing heads appear before index 40, which aligns with PCA’s property that variance decreases as the head index increases. This finding further supports our argument that PCA can approximate preference learning.
@@ 4.3 Test-time Preference Adaptation
The results demonstrate that DRMs achieve the best overall performance across different base models and test sets. The improvement is particularly significant for Gemma-2B-RM, where DRMs improve the single-head baseline from 0.733 to 0.814 on RewardBench and from 0.714 to 0.90 on RPR. Interestingly, the “Shared-Base" ensemble baseline does not outperform the single-head baseline, suggesting that it lacks diversity in its learned reward heads. In contrast, the random-head baseline offers some improvement over the single-head baseline but remains inferior to DRMs. This confirms that DRMs provide a diverse and well-structured set of basis vectors that enable efficient test-time adaptation to user preferences.
@@
Figure 3: Correlation among weight vectors for DRMs. The feature extractor is Gemma-2B-RM.
@@ 4.4 Quantitative Attribute Explainability
Using the weight vectors k for all attributes, we compute their Pearson correlation, providing a quantitative explanation of the dataset’s attributes. The resulting correlation matrix, shown in Figure 3, reveals meaningful relationships between attributes. For instance, “Narrative & Storytelling" strongly correlates with “Humor & Entertainment" and “Linguistic Creativity" (with a correlation of approximately 0.87), which aligns with the idea that humor and creativity enhance storytelling. This correlation suggests that some attributes may be redundant. On the other hand, “Scientific Rigor" negatively correlates with several attributes, including “Chat" and “Narrative & Storytelling" (Pearson $\text{Pearson }r=-0.46$ and $-0.35$, respectively), suggesting that scientific rigor may conflict with other human preferences. Additionally, many attributes show weak or negligible correlations (with absolute Pearson $r<0.1$). Overall, DRMs provide a structured framework for quantitatively explaining attribute relationships, offering deeper insights into benchmark design and multi-attribute evaluation.

### v2

@@ Abstract
Understanding human preferences is crucial for improving foundation models and building personalized AI systems. However, preferences are inherently diverse and complex, making it difficult for traditional reward models to capture their full range. While fine-grained preference data can help, collecting it is expensive and hard to scale. In this paper, we introduce Decomposed Reward Models (DRMs), a novel approach that extracts diverse human preferences from binary comparisons without requiring fine-grained annotations. Our key insight is to represent human preferences as vectors and analyze them using Principal Component Analysis (PCA). By constructing a dataset of embedding differences between preferred and rejected responses, DRMs identify orthogonal basis vectors that capture distinct aspects of preference. These decomposed rewards can be flexibly combined to align with different user needs, offering an interpretable and scalable alternative to traditional reward models. We demonstrate that DRMs effectively extract meaningful preference dimensions (e.g., helpfulness, safety, humor) and adapt to new users without additional training. Our results highlight DRMs as a powerful framework for personalized and interpretable LLM alignment. Our code is available at https://github.com/amandaluof/DRMs
@@
Feng Luo1thanks: Equal contribution., Rui Yang211footnotemark: 1, Hao Sun3, Chunyuan Deng1, Jiarui Yao2, Jingyan Shen4, Huan Zhang2, Hanjie Chen1 1Rice University, 2University of Illinois at Urbana-Champaign, 3University of Cambridge, 4 Columbia University fl38@rice.edu, ry21@illinois.edu, hanjie@rice.edu
@@ 1 Introduction
Reinforcement Learning from Human Feedback (RLHF) Stiennon et al. (2020); Ouyang et al. (2022); Bai et al. (2022); Team et al. (2023); Grattafiori et al. (2024); Liu et al. (2024a) has proven to be a powerful approach for fine-tuning large language models (LLMs) to serve as better general assistants Team et al. (2023); Achiam et al. (2023), AI agents Nakano et al. (2021); Zhao et al. (2024); Yang et al. (2025), and clinical decision support systems Tian et al. (2023); Wang et al. (2024a); Chen et al. (2025). Typically, LLMs are optimized using a scalar reward model trained on human preference data, acting as a proxy for overall user satisfaction. However, this approach has two key limitations: (1) it often reflects the preferences of the majority, potentially marginalizing underrepresented groups Chakraborty et al. (2024); Chidambaram et al. (2024) and failing to capture the full diversity of human preferences, and (2) it struggles to represent the complex, multifaceted, and sometimes conflicting nature of human preferences with a single scalar reward Jang et al. (2023); Rame et al. (2024); Yang et al. (2024c); Zhou et al. (2024).
@@
As demand grows for more personalized LLMs Zhang et al. (2024b), researchers have explored ways to capture fine-grained, multidimensional human preferences. Some studies have introduced datasets that evaluate multiple aspects such as relevance, correctness, completeness, helpfulness, and harmlessness Wu et al. (2023); Wang et al. (2023); Cui et al. (2023); Pitis et al. (2024). Building on these datasets, others have proposed multi-objective optimization methods Qiu et al. (2024) to accommodate diverse user needs Yang et al. (2024c, a); Wang et al. (2024c). However, these approaches face significant challenges: collecting fine-grained human annotations is expensive, and using GPT-generated labels can introduce biases. Considering that large-scale binary preference datasets—where users simply compare two responses—are easier to collect and more widely available, we ask the question: Can we infer multidimensional human preferences directly from large-scale binary comparisons?
@@
To address this question, we propose Decomposed Reward Models (DRMs), a framework that extracts fine-grained human preferences using binary comparison data. Unlike traditional probabilistic models such as Bradley-Terry (BT) Bradley and Terry (1952) or pairwise preference models Jiang et al. (2023), our approach represents human preferences as $d$-dimensional vectors, corresponding to the learned weights of the final linear layer in a reward model. We show that this vector-based representation establishes a connection between Principal Component Analysis (PCA) and preference learning. Our method constructs a dataset of embeddings for all prompt-response pairs and applies PCA to the differences between chosen and rejected response embeddings. This process identifies orthogonal basis vectors, each capturing a distinct human preference direction. These basis vectors can be combined with their corresponding feature extractor to construct reward models. The framework is shown in Figure 1. DRMs are particularly effective for personalized preference learning. Given a small adaptation dataset from a new user, we compute coefficients for the basis vectors and form a linear combination that best aligns with the user’s preferences. This approach allows for flexible, scalable, and interpretable preference modeling without requiring additional training.
@@ 2 Preliminary
The Bradley-Terry (BT) model Bradley and Terry (1952) is a probabilistic framework commonly used in preference learning. Given a prompt $x$ and two responses, $y_{1}$ and $y_{2}$, the BT model defines the probability of $y_{1}$ being preferred over $y_{2}$ as:
@@ 3.1 Vector Representation of Preferences
In this paper, we consider a vector representation of human preference, similar to recent theoretical work Sun et al. (2024); Zhang et al. (2024a); Shen et al. (2025) and empirical studies Sun et al. (2023, 2025). We denote $\phi(x,y)\in\mathbb{R}^{d}$ as a $d$-dimensional feature extractor (e.g., the output from the penultimate layer), and apply a final linear layer $w$ in the reward model, where the reward is computed as $r_{\theta}(x,y)=w^{\top}\phi(x,y)$. Under this formulation, the BT objective can be rewritten as:
@@ 3.3 From Scalar Reward to Diverse Rewards
Our analysis above suggests a connection between the eigenvectors of the covariance matrix and human preferences. A key observation is that the covariance matrix has $d$ eigenvectors—for example, $d=2048$ for gemma-2B Team et al. (2024) and $d=4096$ for Llama-3.1-8B Grattafiori et al. (2024). This means we can extract a large number of meaningful preference vectors $w$ from PCA applied to the embedding dataset $\{z_{i}\}_{i=1}^{N}$.
@@ 3.4 Decomposed Reward Models (DRMs)
(2) Principal Decomposition: Given the dataset $\mathcal{D}_{e}$ of shape $N\times d$, we perform PCA to obtain a set of basis vectors $W=[w_{1},w_{2},\dots,w_{d}]$. Each eigenvector $w_{j}$ captures a distinct directional human preference and can be combined with the feature extractor $\phi(x,y)$ to construct a reward model, leading to a total of $d$ rewards.
To fully utilize DRMs, we aim to represent any human preference as a linear combination of these decomposed rewards. Various methods can achieve this, such as mixture-of-experts Quan (2024); Wang et al. (2024b) and test-time alignment Lee et al. (2024). Since both approaches are fundamentally similar, and test-time alignment can better leverage our method without additional fine-tuning, we adopt it as our implementation. However, DRMs are compatible with any downstream method that requires multiple diverse reward heads.
We implement test-time alignment based on HyRe Lee et al. (2024), which adapts to diverse human preferences by optimizing reward weights given a small adaptation set $\mathcal{D}_{\rm adapt}$. This is practical in real-world scenarios where new users’ data helps refine the strategy to better align with individual preferences. The weights $k_{m}$ are computed based on the loss function over the adaptation set:
$k_{m}=\frac{\exp{(-\mathcal{L}(w_{m},\mathcal{D_{\rm adapt}}))}}{\sum_{j}\exp{(-\mathcal{L}(w_{j},\mathcal{D_{\rm adapt}}))}}$ (7)
@@
$\mathcal{L}(w_{m},\mathcal{D_{\rm adapt}})=-\mathbb{E}_{(x_{i},y_{i}^{c},y_{i}^{r})\sim\mathcal{D}_{\rm adapt}}$ (8)
$\left[\log\sigma\left(w_{m}^{T}(\phi(x_{i},y_{i}^{c})-\phi(x_{i},y_{i}^{r}))\right)\right].$
Here, we also normalize the features based on $\mathcal{D}_{\rm adapt}$ to prevent the loss $\mathcal{L}$ from being dominated by samples with large-scale features. This formulation ensures that weights $k_{m}$ dynamically adjust based on the dataset, assigning higher importance to directions $w_{m}$ with lower loss and reducing the influence of those with higher loss. As a result, we can obtain a soft, weighted sum of DRMs that better aligns with the desired human preference.
@@
Table 1: Performance of top 100 decomposed reward heads. “Single Head” is the trained single-head baseline. “Max Value” refers to the highest score achieved for each attribute, while “Max Head” indicates which specific head attains this maximum score. “Overall” represents the average accuracy of a single head across all attributes.
@@ 4.1 Experimental Setup
Dataset. We choose the mixture2_and_safe_pku dataset [^1] https://huggingface.co/datasets/weqweasdas/preference_dataset_mixture2_and_safe_pku Dong et al. (2024), a collection of 550k pairwise preference samples. This dataset combines diverse data sources, including human-labeled preferences from HH-RLHF Bai et al. (2022) and GPT-labeled preferences from UltraFeedback Cui et al. (2023), making it well-suited for studying diverse preference decomposition. To evaluate the effectiveness of the decomposed reward heads, we test models on two unseen benchmarks with multiple attributes: (1) RewardBench Lambert et al. (2024), a dataset designed to evaluate reward models across various dimensions, including chat quality, safety, and reasoning. (2) Reasonable Preference Reversal (RPR) test set Pitis et al. (2024) focuses on personalized context-aware preference evaluation. From RPR, we sample five fine-grained categories (i.e., User-Friendliness, Narrative Quality, Linguistic Creativity, Scientific Rigor, and Humor), each with over 80 annotated samples. Other categories were excluded due to insufficient data for reliable evaluation.
Base Model. For experiments on decomposed reward heads and test-time adaptation, we use two open-source reward models, Gemma-2B-RM and Llama3-8B-RM Yang et al. (2024b), along with an instruction-tuned language model, gemma-2-9b-it Team et al. (2024), as our base models. To analyze their performance, we keep the backbone fixed as our feature extractors while generating multiple new reward heads for them.
Baselines. We compare the performance of DRMs against several baselines. (1) Single-Head RM fine-tunes a single reward head on top of a fixed backbone using the same dataset as DRMs. (2) Share-Base RM Lee et al. (2024) is a training-based method that employs multiple learnable reward heads while incorporating a frozen prior networkOsband et al. (2023) to maintain diversity, with the final output derived from their combination. (3) Random Head uses multiple reward heads with randomly initialized weights to capture diverse but largely unstructured preferences. Specifically, we experiment with both uniform and Gaussian initialization, followed by L2 normalization to ensure consistency with DRM vectors.
@@
Table 2: Evaluation Results on RPR ($n=5$). We compare DRMs with trained baselines (“Single Head” and “Shared-Base”), and randomly generated multi-head baselines (“Random”). Except for single-head baseline, other methods use HyRe for test-time adaptation. Standard deviation over 20 sampled adaptation sets are reported.
@@
Benchmark Attributes Single Random DRMs
Head (Uniform) (Ours)
@@ 4.2 What information is Captured by DRMs?
We aim to better understand the decomposed reward heads in DRMs. To achieve this, we evaluate the performance of the top 100 reward vectors, ranked by eigenvalue, on both RewardBench and RPR’s fine-grained subsets. Table 1 reports the scores of a trained single-head baseline. The “Max Value” column shows the highest score achieved for each attribute, while the “Max Head” column indicates which reward head achieves this score. We also compare the results with the single-head baseline. The results reveal the following findings:
(1) Diversity and Interpretability: DRMs effectively capture diverse human preferences, with different reward heads excelling at different attributes. For instance, in Gemma-2B-RM, head_9 performs best on “User-Friendliness” (accuracy: 0.798) and “Humor and Entertainment” (0.964), while head_12 excels in “Narrative and Storytelling” (0.825) and “Linguistic Creativity” (0.885). In contrast, the single-head RM fails to capture this diversity, yielding suboptimal performance on attributes such as “User-Friendliness” (0.506) and “Humor and Entertainment” (0.690) for Gemma-2B-RM. These results indicate that DRMs not only capture a broader range of human preferences but also provide interpretable representations that align well with certain preference attributes.
(2) The first head is the most informative: An interesting observation is that the head head_0 consistently achieves the highest overall accuracy for both models. This aligns with expectations, as head_0 corresponds to the eigenvector with the largest variance, i.e., the most informative direction. Furthermore, among the top 100 heads, most of the high-performing heads appear before index 40, which aligns with PCA’s property that the explained variance decreases as the head index increases. This finding further supports our argument that PCA can approximate preference learning.
@@ 4.3 Test-time Preference Adaptation
The results demonstrate that DRMs achieve the best overall performance across different base models and test sets. The improvement is particularly significant for Gemma-2B-RM, where DRMs improve the single-head baseline from 0.733 to 0.814 on RewardBench and from 0.714 to 0.90 on RPR. Interestingly, the “Shared-Base” ensemble baseline does not outperform the single-head baseline, suggesting that it lacks diversity in its learned reward heads. In contrast, the random-head baseline offers some improvement over the single-head baseline but remains inferior to DRMs. This confirms that DRMs provide a diverse and well-structured set of basis vectors that enable efficient test-time adaptation to user preferences.
@@
Figure 3: Correlation among attributes’ weight vectors for DRMs. The feature extractor is Gemma-2B-RM.
@@ 4.4 Quantitative Attribute Explainability
Using the weight vectors k for all attributes, we compute their Pearson correlation, providing a quantitative explanation of the dataset’s attributes. The resulting correlation matrix, shown in Figure 3, reveals meaningful relationships between attributes. For instance, “Narrative & Storytelling” strongly correlates with “Humor & Entertainment” and “Linguistic Creativity” (with a correlation of approximately 0.87), which aligns with the idea that humor and creativity enhance storytelling. This correlation suggests that some attributes may be redundant. On the other hand, “Scientific Rigor” negatively correlates with several attributes, including “Chat” and “Narrative & Storytelling” (Pearson $\text{Pearson }r=-0.46$ and $-0.35$, respectively), suggesting that scientific rigor may conflict with other human preferences. Additionally, many attributes show weak or negligible correlations (with absolute Pearson $r<0.1$). Overall, DRMs provide a structured framework for quantitatively explaining attribute relationships, offering deeper insights into benchmark design and multi-attribute evaluation.